OpenStudy (anonymous):

5 questions please HELP!!!!!!!!!!!!!

OpenStudy (anonymous):

can you help me

OpenStudy (texaschic101):

possibly....what are you having trouble with ?

OpenStudy (anonymous):

ok let me type it for you

OpenStudy (anonymous):

Evaluate (-7) 2 (the 2 has to be at top of the 7 FYI)

OpenStudy (texaschic101):

I do not understand your question...is it just (-7) * 2 or is it -7/2 ??

OpenStudy (texaschic101):

or -7^2 ??

OpenStudy (anonymous):

no its soppose to be like this : (-7) 2 but the two has to be at the top of the 7 but I don't know how to do that

OpenStudy (texaschic101):

you mean -7 squared ??

OpenStudy (anonymous):

yes!!

OpenStudy (texaschic101):

ok.... -7 ^2 = -7 * -7 = 49 or if you had -7^3...it would be -7 * -7 * -7 = 49 * -7 = -343 but be sure and keep track of your signs...neg * neg = pos....neg * pos = neg.

OpenStudy (texaschic101):

does that help ?

OpenStudy (anonymous):

so its -49?

OpenStudy (texaschic101):

not quite.....-7^2 = -7 * -7 = 49.....neg * neg = pos....so you answer is 49, not -49

OpenStudy (anonymous):

ohh thanks can you help me with one more please?

OpenStudy (texaschic101):

sure :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Ok how do you solve this 4 squared do you mulityply 4*2 or do you mulitply 4*4?

OpenStudy (texaschic101):

4^2 = 4 * 4 = 16

OpenStudy (texaschic101):

I am sorry but I got to go...good-luck